A charge experiences a force of 1.5 x 10 -6 N in a uniform electric field of intensity 8.8 N/C. What is the charge?
1
Expert's answer
2020-02-24T11:02:56-0500


A force acts on an electron in an electric field

"F=q\u22c5E"


Where will we write

"q=\frac{F}{E}=\frac{1.5 \cdot 10^{-6}}{8.8}=1.705 \cdot 10^{-7}C"